OpenStudy (anonymous):

Solve the differential equation 4 dy/dx=((e^y)sin^2x)/(ysecx).

OpenStudy (anonymous):

I got the answer e^-y*(-4y-4) = (1/3)*(sin(x))^3 + C However, it does not appear to be correct.

OpenStudy (loser66):

\[4\frac{dy}{dx}= \frac{e^y sin^2x}{y secx}\] right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

\[4\dfrac{y}{e^y} dy = sin^2x cos x dx\] take integral both sides.

OpenStudy (loser66):

\[-4\dfrac{y+1}{e^y}= \frac{sin^3x}{3} + C\]

OpenStudy (anonymous):

That's exactly the answer that I got.

OpenStudy (loser66):

so, why do you make question on your correct answer?

OpenStudy (anonymous):

Did I enter something wrong?

OpenStudy (loser66):

yes, you did. (sin x)^3 not sin (x)^3

OpenStudy (anonymous):

but I also tried sin^3(x) and it did not work.

OpenStudy (loser66):

oh yes, you should put theta, not x

OpenStudy (anonymous):

okay, got it thanks.

OpenStudy (loser66):

ok